# An existence and uniqueness result to evolution equations with sign-changing pseudo-differential operators and its applications to logarithmic Laplacian operators and second-order differential operators without ellipticity

### v1

@@ 3. Fourier transforms and weighted Bessel potential spaces
$(I-\Delta)^{\gamma/2}f:=\mathcal{F}^{-1}\left[\left(1+|\cdot|^{2}\right)\mathcal{F}[f]\right]\in L_{p}(\mathbb{R}^{d}).$
@@
We say that $\gamma_{1}$, $\gamma_{2}$, and $p$ are regularity exponent, weight exponent, and integrability exponent of $H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, respectively.
@@
Same terminologies are used for exponents $\gamma_{1}$, $\gamma_{2}$, and $p$. In other words, $\gamma_{1}$, $\gamma_{2}$, and $p$ are called regularity exponent, weight exponent, and integrability exponent of $H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, respectively.
@@
It is obvious that $(1+|\xi|^{2})^{(p\gamma_{2})/2}$ is in $A_{p}$ since $-d<p\gamma_{2}<d(p-1)$. Moreover, it is also easy to show that $(1+|\xi|^{2})^{(\tilde{\gamma}_{1}-\gamma_{1})/2}$ is a weighted $L_{p}$-multiplier since $\gamma_{1}\geq\tilde{\gamma}_{1}$ (cf. [28]). Therefore, we have
@@
    $\mathcal{F}[f]\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})~{}\text{if and only if}~{}\mathcal{F}^{-1}[f]\in H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d}).$
@@
    $\mathcal{F}[f]\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})~{}\text{if and only if}~{}\mathcal{F}^{-1}[f]\in H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d}).$
@@
Moreover, we omit $in$ and $out$ if $\gamma_{2}=0$ for simpler notation and identification with the unweighted Bessel potential spaces. $w^{p}(t)\mathrm{d}t$ is also skipped if $w(t)\equiv 1$.
@@ 4. Uniqueness of a weak solution with general data
[…]
@@
$\quad+\left\|\frac{1+|\psi(s,\xi)|}{W_{1}(\xi)}\exp\left(\int_{0}^{s}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}s\mathrm{d}\xi\right)}\sup_{s\in(0,t)}\int_{0}^{s}\|\mathcal{F}[f(\rho,\cdot)](\xi)\|_{L_{q}\left(B_{R},W_{2}^{q}(\rho,\xi)\mathrm{d}\xi\right)}\mathrm{d}\rho.$ (5.9)
@@
$\quad+\left(\frac{1}{2\pi}\right)^{\frac{d(2-q)}{2q}}t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1}}_{p,q,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)},$
@@
$\quad+t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,t)\times\mathbb{R}^{d}\right)},$
@@
$\quad+t^{1/p^{\prime}}\left\|\frac{1+|\psi(\rho,\xi)|}{(1+|\xi|^{2})^{\gamma/2}}\exp\left(\int_{0}^{\rho}|\Re[\psi(r,\xi)]|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}\|f\|_{\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,out}\left((0,t)\times\mathbb{R}^{d}\right)},$ (5.13)
@@
$\left\|\frac{1+|a^{ij}(\rho)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)|}{W_{k}(\xi)}\exp\left(\int_{0}^{\rho}\left|\Re[a^{ij}(r)]\xi^{i}\xi^{j}+\Im[b^{j}(r)]\xi^{j}-\Re[c(r)]\right|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}$
@@
$\left\|\frac{1+|a^{ij}(\rho)\xi^{i}\xi^{j}-ib^{j}(t)\xi^{j}-c(t)|}{W_{k}(\xi)}\exp\left(\int_{0}^{\rho}\left|\Re[a^{ij}(r)]\xi^{i}\xi^{j}+\Im[b^{j}(r)]\xi^{j}-\Re[c(r)]\right|\mathrm{d}r\right)\right\|_{L_{p,\infty}\left((0,t)\times B_{R},\mathrm{d}\rho\mathrm{d}\xi\right)}$
@@
×$\qquad\times\exp\left(\int_{0}^{t}\left[R^{2}\sum_{i,j}|a^{ij}(r)|+R\sum_{j}|b^{j}(r)|+|c(r)|\right]\mathrm{d}r\right)$
@@
×$\qquad\times\exp\left(C\left(\sum_{i,j}\|a^{ij}\|_{L_{p}\left((0,t)\right)}R^{2}+\sum_{j}\|b^{j}\|_{L_{p}\left((0,t)\right)}R+\|c\|_{L_{p}\left((0,t)\right)}\right)\right)<\infty,$
@@
Let $p\in[1,\infty]$, $q\in[1,2]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, $u_{0}\in H_{q}^{\gamma_{1}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\gamma_{2}}_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
@@
Let $p\in[1,\infty]$, $q\in(1,2]$, $\gamma_{1},\tilde{\gamma}_{1}\in\left(-\frac{d(p-1)}{p},\infty\right)$, $\gamma_{2},\tilde{\gamma}_{2}\in[0,\infty)$, $u_{0}\in H_{q,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
@@
Let $p\in[1,\infty]$, $q\in(2,\infty]$, $\gamma_{1},\tilde{\gamma}_{1}\in\mathbb{R}$, $\gamma_{2},\tilde{\gamma}_{2}\in\left(\frac{d(q-2)}{2(q-1)},\infty\right)$, $u_{0}\in H_{q,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in

### v2

@@ 3. Fourier transforms and weighted Bessel potential spaces
$(I-\Delta)^{\gamma/2}f:=\mathcal{F}^{-1}\left[\left(1+|\cdot|^{2}\right)^{\gamma/2}\mathcal{F}[f]\right]\in L_{p}(\mathbb{R}^{d}).$
@@
We say that $\gamma_{1}$, $\gamma_{2}$, and $p$ are the regularity exponent, weight exponent, and integrability exponent of $H_{p,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, respectively.
@@
The same terminology is used for exponents $\gamma_{1}$, $\gamma_{2}$, and $p$. In other words, $\gamma_{1}$, $\gamma_{2}$, and $p$ are called the regularity exponent, weight exponent, and integrability exponent of $H_{p,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, respectively.
@@
It is obvious that $(1+|\xi|^{2})^{(p\gamma_{2})/2}$ is in $A_{p}$ since $d<p\gamma_{2}<d(p-1)$. Moreover, it is also easy to show that $(1+|\xi|^{2})^{(\tilde{\gamma}_{1}-\gamma_{1})/2}$ is a weighted $L_{p}$-multiplier since $\gamma_{1}\geq\tilde{\gamma}_{1}$ (cf. [28]). Therefore, we have
@@
[…]
@@
[…]
@@
Moreover, we omit $in$ and outout if $\gamma_{2}=0$ for simpler notation and identification with the unweighted Bessel potential spaces. $w^{p}(t)\mathrm{d}t$ is also skipped if $w(t)\equiv 1$.
@@ 4. Uniqueness of a weak solution with general data
which is described in (4.5). This observation unequivocally implies that a solution $u$ to (1.5) is also unique when viewed as a $\mathcal{F}^{-1}\mathcal{D}^{\prime}(\mathbb{R}^{d})$-valued function defined $(a.e.)$ on $(0,T$).
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
$\qquad\times\exp\left(\int_{0}^{t}\left[R^{2}\sum_{i,j}|a^{ij}(r)|+R\sum_{j}|b^{j}(r)|+|c(r)|\right]\mathrm{d}r\right)$
@@
$\qquad\times\exp\left(C\left(\sum_{i,j}\|a^{ij}\|_{L_{p}\left((0,t)\right)}R^{2}+\sum_{j}\|b^{j}\|_{L_{p}\left((0,t)\right)}R+\|c\|_{L_{p}\left((0,t)\right)}\right)\right)<\infty,$
@@
Let $p\in[1,\infty]$, $q\in[1,2]$, $\gamma_{1},\gamma_{2}\in\mathbb{R}$, $u_{0}\in H_{q}^{\gamma_{1}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\gamma_{2}}_{p,q,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
@@
Let $p\in[1,\infty]$, $q\in(1,2]$, $\gamma_{1},\tilde{\gamma}_{1}\in\left(-\frac{d(p-1)}{p},\infty\right)$, $\gamma_{2},\tilde{\gamma}_{2}\in[0,\infty)$, $u_{0}\in H_{q,in}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,in,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
@@
Let $p\in[1,\infty]$, $q\in(2,\infty]$, $\gamma_{1},\tilde{\gamma}_{1}\in\mathbb{R}$, $\gamma_{2},\tilde{\gamma}_{2}\in\left(\frac{d(q-2)}{2(q-1)},\infty\right)$, $u_{0}\in H_{q,out}^{\gamma_{1},\gamma_{2}}(\mathbb{R}^{d})$, and $f\in\mathbf{H}^{\tilde{\gamma}_{1},\tilde{\gamma}_{2}}_{p,q,out,t\text{-}loc}\left((0,T)\times\mathbb{R}^{d}\right)$. Then there exists a unique weak solution $u$ ((tested by $\mathcal{F}^{-1}\mathcal{D}(\mathbb{R}^{d}))$ to (2.19) in
@@
Declarations of interest
Declarations of interest: none
Data Availability
Data sharing not applicable to this article as no datasets were generated or analysed during the current study.